OpenStudy (anonymous):

Find the zeros: f(x) = 3x^4 - 7x^3 - 6x^2

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

take teh commons, is x^2 then factor out, now u have 3x^4 - 7x^3 - 6x^2 = 0 x^2(3x^2 - 7x - 6) = 0

OpenStudy (raden):

factor out again for term in bracket, it can be x^2 (3x + 2)(x - 3) = 0 now, for the zeroes take each factor and make be zero then solve for x

OpenStudy (anonymous):

thank you :)

OpenStudy (raden):

you're welcome :)

OpenStudy (anonymous):

can you hlep me with another one?

OpenStudy (raden):

sure